OpenStudy (anonymous):

with these equations find the value of x. (x+40) & (3x-10) A.25 B.32.5 C.37.5 D.65

OpenStudy (anonymous):

and here is the picture that goes with it.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

pucture has vertical angles

ganeshie8 (ganeshie8):

your question starts with, "with these equations..." what are the equations ?

ganeshie8 (ganeshie8):

cuz (x+40) & (3x-10) these are not equations, there is no = sign

OpenStudy (anonymous):

(2x+50) & (7x)

ganeshie8 (ganeshie8):

what are the angles ? (2x+50) & (7x) or (x+40) & (3x-10)

ganeshie8 (ganeshie8):

which one you wanto work firs t?

OpenStudy (anonymous):

i wrote the wrong one down for the picture. i meant to ask find angle measurement AEB A.10 B.70 C.110 D.170

ganeshie8 (ganeshie8):

ok, now its making sense :) where is the picture for this ?

OpenStudy (anonymous):

the equations in the picture are (2x+50) and (7x) :)

ganeshie8 (ganeshie8):

vertical angles are equal so 2x+50 = 7x solve x first

ganeshie8 (ganeshie8):

2x+50 = 7x subtract 2x both sides 50 = 5x divide 5 both sides 10 = x x = 10

ganeshie8 (ganeshie8):

AEB = 2x+ 50 put x = 10

ganeshie8 (ganeshie8):

AEB = 2x + 50 = 2(10) + 50 = 20 + 50 = 70